OpenStudy (anonymous):

Solve each logarithmic equation. Express all solutions in exact form. log x^2= (log x)^2

OpenStudy (anonymous):

log(x^2)= (log x)^2 2* log(x) = (log x)^2 2 = log(x) // now i just reverse the two sides log(x) = 2 x = e^2

OpenStudy (anonymous):

the answer is 1100.

OpenStudy (anonymous):

just don't know how???

OpenStudy (anonymous):

well if it is as you typed it above, then x is not 1100. check this out http://www.wolframalpha.com/input/?i=log+x%5E2%3D+%28log+x%29%5E2

OpenStudy (anonymous):

\[\log_{} x ^{2}=(\log_{}x) ^{2}\] that's how the question looks in my book

OpenStudy (anonymous):

one rule of logs is log(x^n) = n*log(x)

OpenStudy (anonymous):

Ok I'll try that. Thank you.

OpenStudy (anonymous):

Tomo, you missed a solution by canceling. You need to factor and set each of the factors equal to zero.

OpenStudy (anonymous):

actually i made a mistake that @Hoa caught. i was thinking in terms of natural log which is why I used e^2. since we are using base 10 we should actually have x = 10^2 = 100

OpenStudy (anonymous):

Wait, so I write it like \[10^{logx ^{2}}=10^{(logx)^{2}}\]

OpenStudy (anonymous):

log(x^2)= (log x)^2 2* log(x) = (log x)^2 2 = log(x) we have log base 10. 10^2 = x

OpenStudy (anonymous):

Thats not corect. You are missing a solution (log(x))^2-2log(x)=0 (log(x))(log(x)-2)=0 log(x)=0 and log(x)=2

OpenStudy (anonymous):

true, that log(x) =0 is missing log(x) = 0 10^0 = x x = 1 AND log(x) = 2 10^2 = x x = 100

OpenStudy (anonymous):

@Xavier it is not "not corect" just miss one solution.

OpenStudy (anonymous):

But why does the answer in the book say {1,100}?

OpenStudy (anonymous):

@joshuaknmcguire that means the solution is 1 and 100

OpenStudy (anonymous):

The procedure there where log(x) is cancelled from both sides is bad.

OpenStudy (anonymous):

sorry, or not and

OpenStudy (anonymous):

OOOOOOOOOOOOH shoot sorry guys!!!! AHH

OpenStudy (anonymous):

aye aye aye I thought it was eleven hundred

OpenStudy (anonymous):

hihihi....

OpenStudy (anonymous):

learn from mistake . good. by the way. Xavier is the best way.

OpenStudy (anonymous):

Thank you @Hoa , @Xavier , and @tomo

OpenStudy (anonymous):

you're welcome